# Local two-dimensional simulations of the ignition of a helium shell detonation on a white dwarf by an impacting stream

### v1

@@ 2 Methods
There are two main components to our model, the CO WD surface with a helium shell and the impacting stream. We model the surface in a plane parallel approximation with the bottom layer representing the core and the top layer representing the helium shell of the WD as seen in Figure 1. We defer the curved surface simulation to future work. The computational configuration for the WD surface is similar to that of Moore et al. (2013) and uses a surface gravity of $4\times 10^{8}$ cm s-2 and a density of $5\times 10^{5}$ g cm-3 at the base of the helium layer (this is varied later in Section 4). This surface gravity approximately corresponds to a 1.0 $\textup{M}_{\odot}$ WD. The base of the helium layer is taken to be $1.5\times 10^{8}$ cm from the bottom of the computational domain. The boundary condition on the bottom of the grid is a variation of a hydrostatic boundary condition (Zingale et al., 2002) where the normal velocity is handled by a reflective boundary condition. We have not observed any issues with handling the normal velocity in such a manner but we will make modifications if they are needed in future simulations. The other boundary conditions will be discussed in the subsections below. The length of the surface in the horizontal direction mimics the circumference of a 1.0 $\textup{M}_{\odot}$ WD. With a radius of $\approx 5\times 10^{8}$ cm, the circumference would be $\approx 30\times 10^{8}$ cm. A surface length of $20\times 10^{8}$ cm was initially used to reduce computational time and, later, the larger surface length of $30\times 10^{8}$ cm was used. A greater surface length is only expected to delay the ignition and hence is not a defining factor. Adaptive mesh refinement is used, with refinement triggered on variations in density, temperature, and helium mass fraction. The largest cell sizes are square and $\approx 310$ km in each dimension, and refinement is allowed to decrease the cell size (by splitting cells in half in each dimension) to a minimum cell size of $\approx 20$ km.
@@ 3.2 Stream study
The formation and propagation of the surface detonation can be seen in Figure 3, which features snapshots in density for the pure helium case. Animated plots of several cases are available from the Zenodo link provided at the end of the manuscript. The phases of evolution are the same for the pure helium and helium with nitrogen cases, though some helium with nitrogen cases ignite earlier as will be discussed below. We demonstrate the pure helium case instead of the other because it is our highest resolution run (cell size $\approx$ 5 km). Table 2 lists all the different runs attempted by varying the velocity for the pure helium case and the helium with nitrogen case. It also lists the peak density at the point and time of ignition. The minimum $\tau_{nuc}/\tau_{dyn}$ just before ignition was found to be $\ll$ 1 (0.05) for the pure helium shell case with a velocity of $6\times 10^{8}$ cm s-1 (lowest resolution).
@@
In addition to the resolution study, one helium with nitrogen shell case was run with a burning limiter (Kushnir et al., 2013), showing ignition as listed in Table 2. While not determinative, this lends support to the robustness of the ignition. Our limiter is described in Boos et al. (2021) and we use a coefficient of 0.1, as done there. The ignition takes place slightly later when the reactions are allowed to be restricted by the limiter, approximately when the next inhomogeneity in the stream interacts with the surface. The site of the ignition in the absence of a limiter still shows the beginnings of a runaway, but the detonation is unable to propagate away from the ignition site, possibly due to the artificial thickening introduced by the limiter. Since the underlying mechanism being discussed here is the continued focusing of the stream, we believe that a slightly later ignition is still a robust demonstration that such focusing can lead to ignition.
@@ 4 Varying the base density of the helium shell
We do not observe a clear threshold for the base helium shell density below which there is no ignition but instead observe shell densities that ignite but cannot propagate a detonation. A base helium shell density of $1.5\times 10^{5}$ g cm-3 and $1.75\times 10^{5}$ g cm-3 ignites the surface but fizzles out soon after. A base helium shell density of $1.9\times 10^{5}$ g cm-3 ignites the surface and a detonation propagates to half the surface length. A base helium shell density of $3\times 10^{5}$ g cm-3 ignites the surface and a detonation propagates around the entire WD. Shen & Moore (2014) have performed a detailed analysis of what shells might host a detonation, so we do not investigate further than finding that shell propagation is the important threshold.
The peak densities at the ignition point and the ignition mode for each of the cases are also listed in Table 3. The peak density seems to slowly decrease with the base density of the helium shell and seem to lie between $1-2\times 10^{6}$ g cm-3. A higher peak density may result in a stronger detonation compared to a lower peak density which would also explain why the low helium shell base density are not able to sustain the detonation. The cases that ignite but are unable to sustain a detonation have weaker shockfronts compared to the cases that ignite and propagate a detonation. This can be seen in Figure 7 which compares four different cases from Table 3 after ignition. Figure 8 demonstrates the dissipation of the detonation for one of the lower base helium shell density cases of $1.9\times 10^{5}$ g cm-3. The shock front separates from the burning front which causes the detonation to fizzle out. Intermediate ignitions (ignition but no propagation) have also been observed in other work (Iwata & Maeda, 2022).
@@
Figure 7: Temperature snapshots of the helium with nitrogen case with different helium shell base densities approximately 3s after the simulation begins (the stream enters the domain). The base density of the model is indicated in the top left hand corner. From Table 3 and from the annotated arrows that point to the shock front in this plot, it is apparent that the strength of the detonation varies with the base densities. The model with $\rho_{\rm{base}}$ = $1.0\times 10^{5}$ g cm-3 does not ignite while $\rho_{\rm{base}}$ = $1.75\times 10^{5}$ g cm-3 and $\rho_{\rm{base}}$ = $1.9\times 10^{5}$ g cm-3 cases ignite the surface, but the detonation propagates to varying distances before fizzling out. The model with $\rho_{\rm{base}}$ = $3.0\times 10^{5}$ g cm-3 has a prominently strong shock front as compared to the other cases and continues to propagate over the entire domain. An animated plot of this evolution is available from the Zenodo link at the end of the manuscript.

### v2

@@ 2 Methods
There are two main components to our model, the CO WD surface with a helium shell and the impacting stream. We model the surface in a plane parallel approximation with the bottom layer representing the core and the top layer representing the helium shell of the WD as seen in Figure 1. We defer the curved surface simulation to future work. The computational configuration for the WD surface is similar to that of Moore et al. (2013) and uses a surface gravity of $4\times 10^{8}$ cm s-2 and a density of $5\times 10^{5}$ g cm-3 at the base of the helium layer (this is varied later in Section 4). This surface gravity approximately corresponds to a 1.0 $\textup{M}_{\odot}$ WD. The scale height, H, is approximately 700 km. The base of the helium layer is taken to be $1.5\times 10^{8}$ cm from the bottom of the computational domain.
The boundary condition on the bottom of the grid is a variation of a hydrostatic boundary condition (Zingale et al., 2002) where the normal velocity is handled by a reflective boundary condition. We have not observed any issues with handling the normal velocity in such a manner but we will make modifications if they are needed in future simulations. The other boundary conditions will be discussed in the subsections below. The length of the surface in the horizontal direction mimics the circumference of a 1.0 $\textup{M}_{\odot}$ WD. With a radius of $\approx 5\times 10^{8}$ cm, the circumference would be $\approx 30\times 10^{8}$ cm. A surface length of $20\times 10^{8}$ cm was initially used to reduce computational time and, later, the larger surface length of $30\times 10^{8}$ cm was used. A greater surface length is only expected to delay the ignition and hence is not a defining factor. Adaptive mesh refinement is used, with refinement triggered on variations in density, temperature, and helium mass fraction. The largest cell sizes are square and $\approx 310$ km in each dimension, and refinement is allowed to decrease the cell size (by splitting cells in half in each dimension) to a minimum cell size of $\approx 20$ km.
@@ 3.2 Stream study
The formation and propagation of the surface detonation can be seen in Figure 3, which features snapshots in density for the pure helium case. Animated plots of several cases are available from the Zenodo link provided at the end of the manuscript. The phases of evolution are the same for the pure helium and helium with nitrogen cases, though some helium with nitrogen cases ignite earlier as will be discussed below. We demonstrate the pure helium case instead of the other because it is our highest resolution run (cell size $\approx$ 5 km). Table 2 lists all the different runs attempted by varying the velocity for the pure helium case and the helium with nitrogen case. It also lists the peak density at the point and time of ignition. The minimum $\tau_{nuc}/\tau_{dyn}$ just before ignition was found to be $\ll$ 1 for the pure helium shell case with a velocity of $6\times 10^{8}$ cm s-1 (lowest resolution) [See appendix]. Therefore, the conditions at hot spots are feasible enough to induce ignition, without the influence of numeric ignition as emphasized by Glasner et al. (2018).
@@
In addition to the resolution study, one helium with nitrogen shell case was run with a burning limiter (Kushnir et al., 2013), showing ignition as listed in Table 2. While not determinative, this lends support to the robustness of the ignition. Our limiter is described in Boos et al. (2021) and we use a value of ${|\Delta\ln\ T|}_{max}$ of 0.1, as done there (Refer to Boos et al. (2021) for a detailed explanation of the burning limiter). The ignition takes place slightly later when the reactions are allowed to be restricted by the limiter, approximately when the next inhomogeneity in the stream interacts with the surface. The site of the ignition in the absence of a limiter still shows the beginnings of a runaway, but the detonation is unable to propagate away from the ignition site, possibly due to the artificial thickening introduced by the limiter. Since the underlying mechanism being discussed here is the continued focusing of the stream, we believe that a slightly later ignition is still a robust demonstration that such focusing can lead to ignition.
@@ 4 Varying the base density of the helium shell
We do not observe a clear threshold for the base helium shell density below which there is no ignition but instead observe shell densities that ignite but cannot propagate a detonation. A base helium shell density of $1.5\times 10^{5}$ g cm-3 and $1.75\times 10^{5}$ g cm-3 ignites the surface but fizzles out soon after. A base helium shell density of $1.9\times 10^{5}$ g cm-3 ignites the surface and a detonation propagates to half the surface length. A base helium shell density of $3\times 10^{5}$ g cm-3 ignites the surface and a detonation propagates around the entire WD. Shen & Moore (2014) have performed a detailed analysis of what shells might host a detonation, that is, use of a larger reaction network and using a more realistic envelope composition reduces the minimum envelope mass that can host a detonation. Therefore, we do not investigate further than finding that shell propagation is the important threshold.
The peak densities at the ignition point and the ignition mode for each of the cases are also listed in Table 3. The peak density seems to slowly decrease with the base density of the helium shell and seems to lie between $1-2\times 10^{6}$ g cm-3. A higher peak density may result in a stronger detonation compared to a lower peak density, which would also explain why the low helium shell base density is not able to sustain the detonation. The cases that ignite but are unable to sustain a detonation have weaker shockfronts compared to the cases that ignite and propagate a detonation. This can be seen in Figure 7 which compares four different cases from Table 3 after ignition. Figure 8 demonstrates the dissipation of the detonation for one of the lower base helium shell density cases of $1.9\times 10^{5}$ g cm-3. The shock front separates from the burning front which causes the detonation to fizzle out. Intermediate ignitions (ignition but no propagation) have also been observed in other work (Iwata & Maeda, 2022).
@@
Figure 7: Temperature snapshots of the helium with nitrogen case with different helium shell base densities approximately 3s after the simulation begins (the stream enters the domain). The base density of the model is indicated in the top left hand corner. From Table 3 and from the annotated arrows that point to the shock front in
this plot, it is apparent that the strength of the detonation varies with the base densities. The model with $\rho_{\rm{base}}$ = $1.0\times 10^{5}$ g cm-3 does not ignite while $\rho_{\rm{base}}$ = $1.75\times 10^{5}$ g cm-3 and $\rho_{\rm{base}}$ = $1.9\times 10^{5}$ g cm-3 cases ignite the surface, but the detonation propagates to varying distances before fizzling out. The model with $\rho_{\rm{base}}$ = $3.0\times 10^{5}$ g cm-3 has a prominently strong shock front as compared to the other cases and continues to propagate over the entire domain. An animated plot of this evolution is available from the Zenodo link at the end of the manuscript.
@@ 5 Summary and Discussion
Appendix A Nuclear Timescale
Figure 9: Evolution of the minimum nuclear timescale in the domain over time. The plot on the right is a zoomed in version of the plot on the left to focus on the minimum nuclear timescale near ignition. The dynamical timescale and the sound crossing time of a cell are also shown in the plot as a pink and purple dashed line respectively. The minimum nuclear timescale seems to drop below the dynamical time scale before ignition ($\tau_{nuc}/\tau_{dyn}\ll 1$) and stays at an approximately constant value soon after the ignition as you would expect it to be for a propagating detonation. The detonation front crosses the whole domain by about 6 s.
When a detonation occurs the nuclear timescale is much smaller than the dynamical timescale (Dan et al., 2012) or $\tau_{nuc}/\tau_{dyn}\ll 1$. $\tau_{nuc}=E_{in}/\dot{E}_{nuc}$ where $E_{in}$ is the internal energy and $\dot{E}_{nuc}$ is the nuclear energy generation rate and $t_{dyn}=H/c_{s}$ where the scale height, $H\approx$ 700 km and the speed of sound, $c_{s}\approx 2.05\times 10^{8}$  cm s-1 in our simulations. In Figure 9 we plotted the minimum $\tau_{nuc}$ over time for one of the pure helium shell cases from Table 2, the case with velocity of the stream = $6.0\times 10^{8}$ cm s-1, surface length = $20\times 10^{8}$ cm, density of stream = $2.5\times 10^{4}$ g cm-3, half width of the stream = $1\times 10^{8}$ cm, stream angle = 35.5, base density of the helium shell = $5\times 10^{5}$ g cm-3 and a cell size $\approx$ 20 km resolution. For technical reasons, this was run with a newer version of our setup having improved boundary conditions and a stream density that gradually decreases from the inner to the outer part of the stream. This results in a similar but not precisely identical ignition, due to minor numerical differences. The plot on the left is the evolution of the minimum $\tau_{nuc}$ over time while the one on the right is the evolution near the point of ignition (at around 4.78 s ). Flash is operator split such that energy deposition and hydrodynamic flow are computed sequentially. This results in $E_{in}$ being different just before the energy deposition rate is computed and just before hydrodynamic flow is computed (after energy deposition). Two curves are plotted, one where $\tau_{nuc}$ was calculated before the nuclear burning and one calculated after the nuclear burning when energy is deposited to $E_{in}$ from the burning. We can see in this figure how the minimum $\tau_{nuc}$ drastically decreases below $\tau_{dyn}$ before ignition and flattens out to a consistent value after ignition as expected for a propagating detonation. The minimum $\tau_{nuc}$ is also lesser than the sound crossing time per cell which is expected since our simulations are still under-resolved. The detonation propagation crosses the entire domain by about 6 s, so that there is no longer an active detonation front after that, and $\tau_{nuc}$ rises back up. Ongoing burning continues in the hot material, but without the detonation front.